1) AN alternating voltage e=100sin314t is applied to a device which offers an ohmic resistance of 50 ohms in one direction while entirely preventing flow of current in opposite direction.Calculate:
a) average and rms value of cuerrent?
b) form factor?


2) An AC was measured as 7mA.Assume the alternating current to be sinusoidal.Calculate: rms,maximum value of current.


3) A voltage is given by v=50 + 30sinx (x=wt; x is = omega * t)in volts. Find the rms value of wave.

i would like to see the detailed steps for solving the problems.... kindly help

Grade:11

2 Answers

Chetan Mandayam Nayakar
12 years ago

1)average=2*(1/pi),rms=2/(2*sqrt(2)),form factor=rms/average

2)

 

Chetan Mandayam Nayakar
12 years ago

2)rms=7*(pi)/sqrt(2),maximum=7*(pi)

3)rms=sqrt(2500+(900/2))=sqrt(2950)